Calculate ΔHf°(SO3) from the following information.

S(s) + O2(g) → SO2(g) ΔHrxn° = -296.8 kJ

SO2(g) + 1/2 O2(g) → SO3(g) ΔHrxn° = -98.9 kJ
Answer #1

Solution: ( -395.7 kJ)

Calculation:

The given reactions are:

S(s) + O2(g) → SO2(g) ΔHrxn° = -296.8 kJ

SO2(g) + 1/2 O2(g) → SO3(g) ΔHrxn° = -98.9 kJ

Adding both equations, we get,

S (s) + 3/2 O2 (g) = SO3(g) , ΔHf° = (-296.8 kJ - 98.9 kJ)

ΔHf° (SO3) =  - 296.8 kJ - 98.9 kJ = - 395.7 kJ